OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = -1/3x + 7 B. y = 3x + 7 C. y = -3x + 7 D. y = 1/3x + 7

OpenStudy (johnweldon1993):

\[\large y = \frac{1}{3}x + 5\] This is the line we have....to find a line perpendicular to it...we first know the the slope of our new line would be the negative reciprocal of this lines slope.. *also known as...when you multiply the slope we get...and the slope we have...you get -1* So Negative = \(\large -\frac{1}{3}\) reciprocal = flip so \(\large -\frac{3}{1} = -3\) This is our new slope...and look we have a point...(2,1) Hmm slope, point? looks like we can use the point slope formula \[\large y - y_o = m(x - x_0)\] replace \(\large m\) with -3 , \(\large y_o\) with 1 , and \(\large x_o\) with 2 \[\large y - 1 = -3(x - 2)\] Distribute this out \[\large y - 1 = -3x + 6\] And add 1 to both sides of the equation \[\large y = -3x + 7\] This would be our equation :)

OpenStudy (anonymous):

What is the equation of the absolute value function?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Do you have choices?